Determine the following present values:

$10,000 received at the end of 13 years; I = 10%

$800 received at the end of each of the next 11 years; I = 14%

$300 received at the beginning of each of the next 4 years I = 12%
Answer #1

ANSWER:

1) I = 10% AND N = 13 YEARS AND FV = 10,000

PV = FV(P/F,I,N)

PV = 10,000(P/F,10%,13)

PV = 10,000 * 0.2897

PV = 2,897

2) I = 14% AND N = 11 YEARS AND ANNUAL PAYMENT RECIEVED = 800

PV = ANNUAL PAYMENT RECIEVED(P/A,I,N)

PV = 800(P/A,14%,11)

PV = 800 * 5.453

PV = 4,362.4

3) I = 12% AND N = 3 YEARS (BECAUSE PAYMENTS ARE MADE AT THE START OF THE YEAR AND HENCE THE FIRST PAYMENT WILL BE IN YEAR 0)  AND ANNUAL PAYMENT RECIEVED = 300

PV = CASH FLOW IN YEAR 0 + ANNUAL PAYMENTS RECIEVED(P/A,I,N)

PV = 300 + 300(P/A,12%,3)

PV = 300 + 300 * 2.402

PV = 300 + 720.6

PV = 1,020.6